OpenStudy (anonymous):

Determine which ordered pair in the form (x, y) is a solution of the given equation. (x + 7)y = 10 (Points : 1) (2, 1) (3, 1) (4, 1) (13, 2)

OpenStudy (anonymous):

hi :D

OpenStudy (anonymous):

Hello!

OpenStudy (anonymous):

does it say \[(x + 7)y = 10\]??

OpenStudy (anonymous):

if so, try \((3,1)\) because \[(3+7)\times 1=10\] is true